A regular octagon is inscribed in a circle of radius 15.8 cm. Find the perimeter of the octagon.
What is the angle of elevation of the sun when a 35-ft mast casts a 20-ft sha

Answer by stanbon(48545)

Sketch the octagon in the circle. It has eight equal sides.
Join the end points of one of the sides to the center of the
circle to form a triangle.
If you did this for each side you would have eight equal
angles at the center sharing 360 Degree.
One of those center angles is 360/8=45 degrees
Therefore the other two angles of the triangle you drew
share 180-45 =135 degrees. Each of them is 135/2=67.5 degrees.
the radii of the circle are 15.8 cm.
Use the Law of Sines to get the following:
15.8/sin 67.5 = x/sin45 where "x" is the length of one side of the octagon.
Using your calculator: x= 11.87
Multiply that by "8" to get the perimeter of the octagon.
2nd problem:
Draw the picture. Then use inverse tangent (35/30) to find the angle
of elevation of the sun.
Cheers,
stan H.